OpenStudy (anonymous):

i need help i have to solve for x 7/x=15/9 but i have to show all my work

OpenStudy (anonymous):

1)cross multiply. 15x = 9*7 15x = 63 2) solve for x x = 63/15

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=7%2Fx+%3D15%2F9